OpenStudy (anonymous):

An equation and the steps to solve it are shown below: 3x - one over two (8x - 2) = 4 Step 1: 3x - 4x + 1 = 4 Step 2: x + 1 = 4 Step 3: x + 1 - 1 = 4 - 1 Step 4: x = 3 Which step was the error made and why?

OpenStudy (anonymous):

@kewlgeek555

OpenStudy (anonymous):

3x -1 /2(8x - 2) = 4 is this the equation mentioned in the question?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

3x (16x - 4) - 1 = 4(16x - 4) will be the first step so It was wrong in the first step itself

OpenStudy (kewlgeek555):

So, let me re-write the original equation: \[3x-0.5(8x-2)=4\]First, you distribute the -0.5 to 8x and -2.\[3x-(-4x)+2=4\]\[3x+4x+2=4\]See how I changed the "-(-" to an addition sign. So I THINK that's where they got wrong, but maybe I'm worng...

OpenStudy (kewlgeek555):

@ganibl it's "-0.5", not "2"

OpenStudy (anonymous):

@kewlgeek555 Does 1/2 and 0.5 are not same :o

OpenStudy (anonymous):

*facepalm*

OpenStudy (kewlgeek555):

I think it does @ganibl

OpenStudy (alekos):

babydoll step 1 is right step 2 is wrong should be -x +1 = 4

OpenStudy (alekos):

then we have -x = 3

OpenStudy (alekos):

so x = -3